Hi,
It has been many years since I've done any Math, and I forgot how to do some simple stuff. Could you please help me to do the steps for this problem?:
1/4=P(A)/(1-P(A)) - - I multiply both sides by the denominator to isolate P(A)
(0.25)(1-P(A))=P(A) - - I multiply both sides by 4/4 to convert to a decimal
0.25-(0.25)P(A)=P(A) - - distribute
0.25=(1.25)P(A) - - ?
P(A)=0.25/1.25 - - ??
P(A)=0.20 .
Thank you very much for your help.
Stephanie

Answer by Alan3354(69443)

1/4=P(A)/(1-P(A)) - - I multiply both sides by the denominator to isolate P(A)
(0.25)(1-P(A))=P(A) - - I multiply both sides by 4/4 to convert to a decimal
0.25-(0.25)P(A)=P(A) - - distribute
0.25=(1.25)P(A) - - ?
P(A)=0.25/1.25 - - ??
P(A)=0.20 .
===================
1/4 = P/(1-P)
Cross multiply (a shortcut and eliminates fractions)
1-P = 4P
Add P
1 = 5P
P = 1/5 = 0.2